# Accurate estimation of feature importance faithfulness for tree models

### v1

@@ 4.3 Comparing squared prediction gap algorithms
The comparison between the algorithms was carried out in different settings. The variable parameters were:
  • Model type $m$ trained on a selected dataset. In our case, there were four models in total (Section 4.2) trained on two datasets (Section 4.1).
  • The standard deviation $\sigma$ of a Gaussian used to perturb a feature. In our experiment, the evaluated values were $\{0.01,0,03,0.1,0.3\}$.
[…]
$\text{NMAE}(y,\hat{y})=\frac{\sum_{j=1}^{N}\mid y_{j}-\hat{y_{j}}\mid}{\sum_{j=1}^{N}{\mid y_{j}\mid}}$
where $y_{j}$ is the correct value of the prediction gap calculated by the exact algorithm, $\hat{y_{j}}$ is the prediction gap calculated by the sampling algorithm. NMAE was used as opposed to the relative error to suppress the effect of relatively large floating-point errors present when operating on very small numbers.
@@
The results obtained in the experiment – the error as a function of the number of iterations – are depicted using the “Mean” curve for $\sigma=0.3$ and Bigger model for the Red Wine Quality dataset in Figure 1 and for $\sigma=0.03$ and Bigger model for the Californian Housing dataset in Figure 2. Analogous figures for other settings $(m,\sigma)$ can be found in Appendix B.
Additionally, the figures contain separate error curves for some selected sizes $|S|$ of the perturbed features set. As can be observed for a given iteration, the NMAE error behaves similarly for all subset sizes. Overall, the difference between the exact and sampling algorithms decreases as the number of iterations in the sampling algorithm increases. This is indeed anticipated since Monte Carlo sampling yields better accuracy with iteration increase; in fact the relative error decreases proportionally to $1/\sqrt{i}$.
Figure 1: Results of precision comparison described in 4.3 for $\sigma=0.3$ and for Bigger model for Red Wine Quality dataset.
Figure 2: Results of precision comparison described in 4.3 for $\sigma=0.03$ and for Bigger model for California housing dataset.
For the fair comparison, we chose the number of iterations (out of $9$ considered as explained above) in the sampling algorithm that came closest to the running time of the exact algorithm for each setting. The chosen number of iterations and the associated NMAE are shown in Tables 1, 2, 3 4.
Table 1: Number of iterations needed to perform to match execution time of exact algorithm on single tree model for Wine Quality Dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 100 0.08
0.03 100 0.14
0.1 100 0.15
0.3 1000 0.05
Table 2: Number of iterations needed to perform to match execution time of exact algorithm on Bigger model for Wine Quality Dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 100 0.08
0.03 100 0.11
0.1 100 0.12
0.3 1000 0.04
Table 3: Number of iterations needed to perform to match execution time of exact algorithm on single tree model for Californian Housing dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 100 0.13
0.03 100 0.15
0.1 100 0.15
0.3 1000 0.04
Table 4: Number of iterations needed to perform to match execution time of exact algorithm on Bigger tree model for Californian Housing dataset described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 500 0.3
0.03 500 0.14
0.1 1000 0.06
0.3 2000 0.03
We conclude that the “fair” comparison confirms that the Monte Carlo simulation gives noticeably less accurate results than our exact method. We observe that the error generally decreases when increasing the size of the model. This is justified by the fact that our algorithm’s running time increases quadratically with $n$, whereas a single iteration of the Monte Carlo method runs in time at most linear in $n$. As a result, the bigger the model size is, the more iterations the Monte Carlo method can perform within the same time budget. Moreover, the error decreases with the standard deviation of the distribution used to obtain perturbations. This is also expected: for smaller perturbations, the floating-point rounding errors are of higher significance.
Table 5: Number of iterations needed to perform to match execution time of exact algorithm on Bigger Californian Housing model described in 4.3.
$\sigma$ Number of iterations Relative error
0.01 500 0.30
0.03 500 0.14
0.1 1000 0.06
0.3 2000 0.03
5 Feature ranking using $\mathrm{PG}^{2}$
We propose a method for ranking the features of a model using the squared prediction gap called greedy $\text{PG}^{2}$. Let $x\in\mathbb{R}^{d}$. Suppose the feature perturbations $\delta_{i}$ all come from $\mathcal{D}=\mathcal{N}(0,\sigma)$. The ranking $\pi_{\mathrm{PG}}^{\sigma}(x)$ is constructed in the following inductive manner.
  1. 1.
    For every $i\in[d]$, compute $p_{i}=\mathrm{PG}^{2}(x,\{i\})$. Set the most important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(1)$ to be the feature $i$ for which $p_{i}$ is the highest.
  2. 2.
    Assume $\pi^{\sigma}_{\mathrm{PG}}(x)(1),...,\pi^{\sigma}_{\mathrm{PG}}(x)(l)$ are already computed. Let ${S=\{\pi^{\sigma}_{\mathrm{PG}}(x)(1),...,\pi^{\sigma}_{\mathrm{PG}}(x)(l)\}}$. For every $i\in[d]\setminus S$, compute $p_{i}:=\mathrm{PG}^{2}(x,S\cup\{i\})$. Let the next important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(l+1)$ be the feature $i$ for which $p_{i}$ is the highest.
In the following, we compare the feature rankings obtained using the above method with the rankings obtained using the popular SHAP attribution framework [10].
For a given $x\in\mathbb{R}^{d}$, SHAP produces an attribution vector $(\phi)_{i=1}^{d}$ whose elements sum up to $f(x)-\mathbb{E}[f(x)]$ but can be both positive and negative [12]. We define the ranking $\pi_{\text{SHAP}}(x)$ to be obtained by sorting the features $[d]$ by the absolute value of $\phi_{i}$, from highest to lowest. This is justified by the fact that the predictions $f(x)$ can be smaller than the expected value; in such a case, all the attributions $\phi_{i}$ might be negative.
5.1 Global comparison vs SHAP wrt. $\mathrm{PGI}^{2}$
We perform a global comparison of $\pi^{\sigma}_{\mathrm{PG}}$ and $\pi_{\text{SHAP}}$ wrt. the $\mathrm{PGI}^{2}$ metric, that is, we check which ranking method gives higher $\mathrm{PGI}^{2}$ scores on average. More formally, fix a dataset/model combination $m$ and a perturbation distribution $\mathcal{D}^{\prime}=\mathcal{N}(0,\sigma^{\prime})$, where potentially $\sigma^{\prime}\neq\sigma$, as we did in Section 4.3. Let $X\subseteq{\mathbb{R}}^{d}$ be the corresponding dataset used in $m$. Then, define the average $\mathrm{PGI}^{2}$ scores as
$\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi_{\text{SHAP}}(x)),$
$\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi_{\mathrm{PGI}}^{\sigma}(x)).$
Figures 345 and 6 present the comparison of the average scores $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ and $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ for $\sigma=0.3$ as a function of $\sigma^{\prime}$. We observe that even if the parameter $\sigma$ used when computing the ranking $\pi_{\mathrm{PG}}^{\sigma}$ differs from the parameter $\sigma^{\prime}$ used for computing the $\mathrm{PGI}^{2}$ metrics, $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ is consistently better than $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ for the Bigger models. For Single tree models, this is the case only for sufficiently large $\sigma^{\prime}$, but in particular for $\sigma^{\prime}=\sigma$. The figures presenting the results for other choices of $\sigma$ can be found in Appendix B.2.
Figure 3: Single tree model for the Wine Quality dataset.
@@ 5 Feature ranking using $\mathrm{PG}^{2}$
Figure 4: Bigger model for the Wine Quality dataset.
Figure 5: Single tree model for the Housing dataset.
Figure 6: Bigger model for the Housing dataset
@@
We conclude that the greedy optimization of the $\mathrm{PGI}^{2}(x,\pi)$ metric performs reasonably well, even though its correctness could only be guaranteed if it optimized the $\mathrm{PGI}^{2}(x,\pi,1)$ metric (see Section 2) instead.
5.2 Conciseness comparison vs SHAP
We also compare the global conciseness of the feature rankings $\pi_{\text{SHAP}}$ and $\pi_{\mathrm{PG}}^{\sigma}$. Intuitively, we would like to declare a feature ranking method more concise if it tends to produce rankings that identify the most important features of the model more clearly. For example, the information that two or three features of the model are identified as much more important than the others seems to say more about the model’s behavior in general than saying that all the features are equally important on average.
Our conciseness evaluation is performed using four different methods of converting a feature ranking $\pi(x)$ to a score vector $r^{\pi}(x):[d]\to\mathbb{R}_{\geq 0}$: geom, top1, top2, top3:
  • geom sets $(r^{\pi}(x))_{i}=1/2^{k}$ if $\pi(k)=i$, that is, the scores increase geometrically with the importance.
[…]
Having fixed (as before) the model-dataset combination $m$, $\sigma$, and a scoring method, we compute the aggregate feature scores $R_{i}$ through all data points $x$ from the test dataset $X$:
$R_{i}=\sum_{x\in X}(r^{\pi}(x))_{i}.$
Finally, the conciseness $C_{\pi}$ is computed as the entropy of the normalized vector $R_{1},\ldots,R_{d}$:
$C_{\pi}=\sum_{i=1}^{d}\frac{R_{i}}{\sum_{j}R_{j}}\log\left(\frac{R_{i}}{\sum_{j}R_{j}}\right).$
Intuitively, higher entropy means more uncertainty in the results.
Tables 6,78, and 9 present the results of our comparison for $\sigma\in\{0.1,0.3\}$ for Wine Quality models, and additionally for $\sigma=1$ for the Housing models. We observe that for the “single tree” models, the entropy that SHAP ranking yields is slightly smaller, for most of the scoring methods. However, there is not much uncertainty in the produced rankings since the models base their predictions only on a few features. Indeed, in the case of “single tree” models for both Wine Quality and Californian Housing datasets, only three features were present in trees.
For the bigger models consisting of numerous trees, all the features are used. In this more complex case, where the top features are not identified immediately, the greedy $\mathrm{PG}$-based ranking yields more concise rankings for the geom, top2 and top3 metrics for the Wine Quality dataset. For the Housing dataset, however, the obtained entropies for $\pi_{\text{SHAP}}$ are also consistently smaller than for the $\pi_{\mathrm{PG}}^{\sigma}$ rankings for $\sigma\in\{0.1,0.3\}$. Only for $\sigma=1$, the greedy $\mathrm{PG}^{2}$ method looks generally more concise.
Table 6: Entropy comparison between SHAP rankings and PG rankings for Bigger model for Wine Quality Dataset.
score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
geom 3.245 3.285 3.157 2.740
top1 2.643 3.059 2.792 1.730
top2 3.120 3.232 3.019 2.402
top3 3.305 3.300 3.189 2.846
Table 7: Entropy comparison between SHAP rankings and PG rankings for Bigger model for Californian Housing Dataset.
score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
geom 2.570 2.774 2.624 2.383
top1 1.748 2.324 1.979 0.534
top2 2.394 2.651 2.377 2.176
top3 2.634 2.834 2.676 2.622
Table 8: Entropy comparison between SHAP rankings and PG rankings for Single tree model for Wine Quality Dataset.
score SHAP PG, $\sigma=0.1$ PG, $\sigma=0.3$ PG, $\sigma=1.0$
geom 2.644 2.463 2.328 2.316
top1 1.725 1.482 1.372 1.305
top2 2.112 1.988 1.654 1.585
top3 2.432 2.207 2.073 2.012
Table 9: Entropy comparison between SHAP rankings and PG rankings for Single tree model for Californian Housing Dataset.
score SHAP PG $\sigma=0.1$ PG $\sigma=0.3$ PG $\sigma=1.0$
geom 2.094 2.197 2.175 2.138
top1 1.077 0.884 0.722 0.031
top2 1.549 1.580 1.602 1.633
top3 1.585 1.998 2.020 2.066

### v3

@@ Introduction
Results
Accuracy
To estimate the accuracy of sampling-based methods as a function of the iteration count, we computed the Normalised Mean Absolute Error (NMAE), defined as:
$\text{NMAE}(y,\hat{y})=\frac{\sum_{j=1}^{N}\mid y_{j}-\hat{y_{j}}\mid}{\sum_{j=1}^{N}{\mid y_{j}\mid}},$
where $y_{j}$ is the respective value $\mathrm{PG}(x_{j},S_{j})$ computed by the exact algorithm, and $\hat{y_{j}}$ is the respective estimate obtained by the sampling algorithm. As opposed to the more popular MAPE error, NMAE was used because it is well defined in cases when true prediction is 00, which is often the case (for example when the perturbation is small).
@@ Results
The results obtained in the experiment – the error as a function of the number of iterations – are depicted using the “Mean Monte Carlo” and “Mean Quasi Monte Carlo” curves for $\sigma=0.3$ and Bigger model for the Red Wine Quality dataset in Figure 1. Analogous figures for other datasets, models and $(m,\sigma)$ can be found in the Appendix.
@@
Figure 1: NMAE for $\sigma=0.3$ and for Bigger model for Red Wine Quality dataset.
Overall, the difference between the exact and sampling algorithms decreases as the number of iterations in the sampling algorithm increases. This is indeed anticipated since MC sampling yields better accuracy with iteration increase; in fact, the relative error decreases proportionally to $1/\sqrt{i}$. Additionally, the error for the QMC method is significantly lower, especially for the lower count of iterations. Moreover, the consistent $\text{NMAE}(y,\hat{y})$ error drop with the increase of the iteration count of the sampling methods indicates that our closed-form algorithm is numerically stable.
Efficiency
The running time of the sampling methods depends, clearly, linearly on the iterations count. As a result, the decision about which method to choose should depend on the acceptable accuracy. To this end, for each of the considered settings (dataset and parameters), we identified the number of iterations (out of $13$ considered thresholds as explained above) in the respective sampling algorithm that came closest to the running time of the exact algorithm for each setting. These corresponding numbers of iterations and the associated NMAEs are shown in Table 1. A larger acceptable NMAE than the corresponding threshold value suggests using the respective sampling-based method, and for smaller NMAEs than that, the exact algorithm is preferable.
Monte Carlo Quasi Monte Carlo
$\sigma$ Iterations NMAE Iterations NMAE
Single tree Red Wine Quality model
0.1 100 0.132 100 0.04
0.3 100 0.136 100 0.049
1.0 100 0.122 100 0.054
Bigger Red Wine Quality model
0.1 4000 0.019 4000 0.003
0.3 8000 0.014 8000 0.002
1.0 15000 0.009 15000 0.001
Single California Housing model
0.1 100 0.158 100 0.051
0.3 100 0.14 100 0.043
1.0 100 0.129 100 0.049
Bigger California Housing model
0.1 4000 0.018 4000 0.002
0.3 10000 0.011 10000 0.001
1.0 15000 0.01 15000 0.001
Single Parkinson Telemonitoring model
0.1 500 0.071 100 0.037
0.3 500 0.064 500 0.011
1.0 500 0.085 100 0.092
Bigger Parkinson Telemonitoring model
0.1 6000 0.015 4000 0.004
0.3 15000 0.009 15000 0.002
1.0 20000 0.008 20000 0.002
Table 1: Iterations needed to match the execution time of our exact algorithm, and the corresponding NMAEs.
We observe that the vanilla Monte Carlo simulation gives noticeably less accurate results than our exact method with a comparable allocated time budget. We observe that the error generally decreases when increasing the size of the model. This is justified by the fact that our algorithm’s running time increases quadratically with $n$, whereas a single iteration of the Monte Carlo and Quasi Monte Carlo method runs in time at most linear in $n$. As a result, the bigger the model size is, the more iterations the Monte Carlo and Quasi Monte Carlo methods can perform within the same time budget. Moreover, for Monte Carlo the error decreases with the standard deviation of the distribution used to obtain perturbations.
Feature Ranking Using $\mathrm{PG}^{2}$
We also study and evaluate a method for ranking the features of a model using the squared prediction gap called greedy $\text{PG}^{2}$. Let $x\in\mathbb{R}^{d}$. Suppose the feature perturbations $\delta_{i}$ all come from $\mathcal{D}=\mathcal{N}(0,\sigma^{2})$. The ranking $\pi_{\mathrm{PG}}^{\sigma}(x)$ is constructed in the following inductive manner.
For $l=0,1,\ldots,d-1$, assume we have already computed the $l$ top features $\pi^{\sigma}_{\mathrm{PG}}(x)(1),\ldots,\pi^{\sigma}_{\mathrm{PG}}(x)(l)$. Let ${S=\{\pi^{\sigma}_{\mathrm{PG}}(x)(1),...,\pi^{\sigma}_{\mathrm{PG}}(x)(l)\}}$. For every $i\in[d]\setminus S$, compute $p_{i}:=\mathrm{PG}^{2}(x,S\cup\{i\})$. We define the next important feature $\pi^{\sigma}_{\mathrm{PG}}(x)(l+1)$ to be the feature $i$ for which $p_{i}$ is the highest.
In the following, we compare the feature rankings obtained using the above method with the rankings obtained using the popular SHAP attribution framework (Lundberg et al. 2020) for tree ensembles, concretely using the tree_path_dependent feature perturbation method.
For a given $x\in\mathbb{R}^{d}$, SHAP produces an attribution vector $(\phi)_{i=1}^{d}$ whose elements sum up to $f(x)-\mathbb{E}[f(x)]$ but can be both positive and negative (Lundberg and Lee 2017). We define the ranking $\pi_{\text{SHAP}}(x)$ to be obtained by sorting the features $[d]$ by the absolute value of $\phi_{i}$, from highest to lowest. This is justified by the fact that the predictions $f(x)$ can be smaller than the expected value; in such a case, all the attributions $\phi_{i}$ might be negative.
A natural question arises whether this ranking method produces rankings similar or dissimilar to SHAP rankings. Table 2 shows the ratio of datapoints for which the top $k$ most important features are the same for $\pi^{\sigma}_{\mathrm{PG}}$ and $\pi_{\text{SHAP}}$. As we see, this ratio usually decreases with $k$, increases with $\sigma$, and is lower for the bigger models than for the single tree models. However, even the rankings for single tree models pick the same most important feature as SHAP less than 76% of the time. In general, the rankings produced are quite different, so it is worthwhile to study these differences and perform a comparative analysis of the rankings.
$\sigma$ $k=1$ $k=2$ $k=3$
Single tree Red Wine Quality model
0.1 0.328 0.259 0.031
0.3 0.412 0.591 0.428
1.0 0.419 0.613 0.506
Bigger Red Wine Quality model
0.1 0.131 0.041 0.022
0.3 0.212 0.094 0.056
1.0 0.391 0.181 0.103
Single California Housing model
0.1 0.501 0.263 0.031
0.3 0.599 0.569 0.349
1.0 0.712 0.471 0.504
Bigger California Housing model
0.1 0.194 0.08 0.052
0.3 0.3 0.122 0.056
1.0 0.569 0.243 0.08
Single Parkinson Telemonitoring model
0.1 0.402 0.143 0.066
0.3 0.569 0.32 0.237
1.0 0.759 0.414 0.22
Bigger Parkinson Telemonitoring model
0.1 0.137 0.028 0.007
0.3 0.225 0.061 0.009
1.0 0.368 0.114 0.031
Table 2: Ratio of datapoints having the same $k$ most important features in $\pi_{\mathrm{PG}}^{\sigma}$ as in $\pi_{\text{SHAP}}$
Comparison vs SHAP wrt. $\mathrm{PGI}^{2}$
We perform a global comparison of $\pi^{\sigma}_{\mathrm{PG}}$ and $\pi_{\text{SHAP}}$ wrt. the $\mathrm{PGI}^{2}$ metric, that is, we check which ranking method gives higher $\mathrm{PGI}^{2}$ scores on average. More formally, fix a dataset/model combination $m$ and a perturbation distribution $\mathcal{D}^{\prime}=\mathcal{N}(0,(\sigma^{\prime})^{2})$, where potentially $\sigma^{\prime}\neq\sigma$, for perturbations used in computing $\mathrm{PG}^{2}(x,\cdot)$ in (3). Let $X\subseteq{\mathbb{R}}^{d}$ be the corresponding dataset used in $m$. Then, for $\pi\in\{\pi_{\text{SHAP}},\pi_{\mathrm{PGI}}^{\sigma}\}$, define the average $\mathrm{PGI}^{2}$ scores as
$\overline{{\mathrm{PGI}}^{2}}(\pi)$ $=\frac{1}{|X|}\sum_{x\in X}\mathrm{PGI}^{2}(x,\pi(x)).$
Table 3 presents the average scores $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ and $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ for $\sigma\in\{0.1,0.3,1.0\}$ as a function of $\sigma^{\prime}$ for Bigger models (an analogous table for Single models can be found in the Appendix). We observe that even if the parameter $\sigma$ used when computing the ranking $\pi_{\mathrm{PG}}^{\sigma}$ differs from the parameter $\sigma^{\prime}$ used for computing the $\mathrm{PGI}^{2}$ metrics, $\overline{{\mathrm{PGI}}^{2}}(\pi_{\mathrm{PGI}}^{\sigma})$ is almost always better than $\overline{{\mathrm{PGI}}^{2}}(\pi_{\text{SHAP}})$ for the Bigger models, with a slight exception in the ${\sigma^{\prime}=1.0,\sigma=0.1}$ pair case, when $\sigma,\sigma^{\prime}$ differ a lot.
$\sigma^{\prime}$ SHAP $\sigma$ = $0.1$ $\sigma$ = $0.3$ $\sigma$ = $1.0$
Bigger PGI Red Wine Quality model
0.1 0.013 0.024 0.022 0.018
0.3 0.051 0.068 0.078 0.067
1.0 0.221 0.192 0.266 0.302
Bigger PGI California Housing model
0.1 3.346e+08 4.205e+08 3.841e+08 3.065e+08
0.3 1.209e+09 1.297e+09 1.374e+09 1.259e+09
1.0 7.807e+09 6.963e+09 8.266e+09 8.728e+09
Bigger PGI Parkinson Telemonitoring model
0.1 2.743 4.425 4.015 3.199
0.3 6.341 9.458 11.878 10.143
1.0 16.298 19.56 30.188 34.013
Table 3: Greedy $\mathrm{PG}^{2}$ vs SHAP rankings wrt. $\mathrm{PGI}^{2}$.
@@
We conclude that the greedy optimization of the $\mathrm{PGI}^{2}(x,\pi)$ metric performs reasonably well, even though its correctness could only be guaranteed if it optimized the $\mathrm{PGI}(x,\{\pi_{1}\})$ metric (where only the single most important feature is perturbed) instead of (3).
Comparison vs SHAP with Feature Removing
We compare the rankings $\pi_{\text{SHAP}}$ and $\pi_{\mathrm{PG}}^{\sigma}$ using feature removing. Intuitively, removing information about the value of a highly important feature should result in greater model error than removing information about the value of a less important feature. Thus, we can assess a ranking by measuring the error of the model when the features the ranking deems most important are removed. More precisely, for a ranking $\pi$ and for each $k=1,2,...$ we measure the RMSE of $\xi(x,[d]\setminus\pi(x)[1..k])$ over all $x$ from the test dataset, where $\xi(\cdot,S)$ is a function giving model predictions based only on features from $S$. We have two such functions $\xi$ (and thus two different methods of comparing rankings) defined below.
The first method is feature randomising. We define $\xi_{random}(x,S):=\mathbb{E}[m(X)]$ where $m$ is the model and $X$ is a random variable with $X_{i}\equiv x_{i}$ for $i\in S$ and $X_{j}$ has the distribution of the $j$-th feature over the test set for $j\not\in S$. In practice, the expected value is approximated by sampling $m(X)$ 100 times (where each feature is sampled independently) and taking the average.
The other method follows the remove-and-retrain approach described in (Hooker et al. 2019). For $k=1,2,3,...$ and for each subset of features $S\subset[d]$ such that $|S|=d-k$ we train a XGBoost model where we only use the features in $S$. That is, the $k$ features that are not in $S$ are not available for the training of the model. This way, we obtain a family of models $(m_{S})_{S\subset[d]}$ where each model $m_{S}$ gives predictions based only on values of features in $S$. We define $\xi_{retrain}(x,S):=m_{S}(x)$.
The comparison was performed for $k=1,2,3$. The results are shown in Tables in the Appendix. The results for the retraining method suggest that $\pi_{\text{SHAP}}$ picks the most important features more accurately than $\pi_{\mathrm{PG}}^{\sigma}$. The results for the randomising method show no significant difference between $\pi_{\text{SHAP}}$ and $\pi_{\mathrm{PG}}^{\sigma}$ for Red Wine Quality and Parkinson Telemonitoring models. However, for the Housing model we see that $\pi_{\mathrm{PG}}^{\sigma}$ with $\sigma\in\{0.1,0.3\}$ give better results than $\pi_{\text{SHAP}}$. In general, this comparison doesn’t give a simple conclusion that one ranking method is clearly better than the other. Each of them can show advantage over the other depending on the dataset and assessment method.
Conclusion
In this work, we proposed a cleanly defined faithfulness metric $\mathrm{PGI}^{2}$ that can be computed exactly in polynomial time on tree ensemble models. We designed a quadratic algorithm for that. Our experimental evaluation showed that the algorithm is numerically stable and superior to Monte Carlo-based methods if very accurate results are desired.
We also proposed a natural feature ranking method inspired by $\mathrm{PGI}^{2}$ optimization. The method generally identifies different most important features than the state-of-the-art SHAP explainer while offering similar performance wrt. different faithfulness metrics. The $\mathrm{PG}^{2}$-based ranking may thus offer a viable alternative for assessing feature importance in tree-based models.
Acknowledgments
Partially supported by the ERC PoC grant EXALT no 101082299 and the National Science Centre (NCN) grant no. 2020/37/B/ST6/04179.